Galvanized nails are iron nails that have been plated with zinc to prevent rusting. The relevant reaction is

Zn2+(aq)+2e−→Zn(s)

For a large batch of nails, a manufacturer needs to plate a total zinc mass of 3.20 kg on the surface to get adequate coverage.

Part C:

With a power supply able to produce 125 A of current, how long must the nails stay in the electrolysis chamber to receive 9.44×106 C of charge?

Express your answer in hours.
Answer #1

C)
I = 125 A
Q = 9.44*10^6 C

Use:
I = Q/t
125 = (9.44*10^6)/t
t = 75520 s
= 75520/3600 hr
= 21.0 hr

Answer: 21.0